# 5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8

### v2

@@ 4.1 Proposed TLL-MMSE in rank-deficient channel
$+\sum_{j=1}^{K}\sigma^{2}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}+\sigma^{2}_{z}\mathbf{I}_{k})\mathbf{U}_{k}$
@@
$+\sum_{j=1}^{K}\sigma^{2}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}$
@@ 4.2 Proposed NSP-MTP-MRP in LoS channel
$+\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{n}+\mathbf{u}_{0}^{H}\mathbf{z},$ (63)
where $\mathbf{v}_{0}$ and $\mathbf{u}_{0}^{H}$ are transmit and receive beamforming vectors of user $1$ from reflect channel respectively, and $\mathbf{\Theta}$ is the reflection coefficient matrix of IRS. And $\mathbf{v}_{k}$ and $\mathbf{u}_{k}^{H}$ are transmit and receive beamforming vectors of user $k$ from direct channel.
@@
$+\sigma^{2}\|\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\|^{2}+\sigma^{2}_{z}\|\mathbf{u}_{0}^{H}\|^{2}.$ (64)
@@
where $A_{0}=\sigma^{2}_{z}\|\mathbf{u}_{0}^{H}\|^{2}+\sigma^{2}\|\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\|^{2}$.
@@
$+\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{n}+\mathbf{u}_{k}^{H}\mathbf{z}.$ (67)
@@
where $A_{k}=\sigma^{2}\|\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\|^{2}+\sigma^{2}_{z}\|\mathbf{u}_{k}^{H}\|^{2}$.
@@
$P_{I}$ $=\hat{\mu}^{2}(\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+\sigma^{2}),$ (82)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{I}}{\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+\sigma^{2}}}.$ (83)
@@ 4.3 Proposed SO-MMSE in LoS channel
$P_{I}$ $=\hat{\mu}^{2}(\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+\sigma^{2}),$ (85)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{I}}{\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+\sigma^{2}}}.$ (86)
@@
$+\sigma^{2}\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k}\mathbf{u}_{k}+\sigma^{2}_{z}\mathbf{u}_{k}^{H}\mathbf{u}_{k},$ (88)
@@
$\frac{\partial f(\mathbf{u}_{k})}{\partial\mathbf{u}_{k}}$ $=2(\sigma^{2}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k}+\sigma^{2}_{z}\mathbf{I}_{k}+\mathbf{H}_{dk}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{H}_{dk}^{H}$
@@
$\mathbf{H}_{dk}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{H}_{dk}^{H}+\sigma^{2}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k})^{-1}(\mathbf{H}_{dk}\mathbf{v}_{k}).$ (92)
@@
$\mathbf{H}_{d1}\mathbf{v}_{1}\mathbf{v}_{1}^{H}\mathbf{H}_{d1}^{H}+\sigma^{2}\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{1})^{-1}(\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{v}_{0}).$ (93)
@@ 5 Simulation Results
In the following, simulation results are presented to verify the DoF gains achieved by IRS and assessed the corresponding sum-rate performance of the designed three methods. Below, simulations are conducted in two different situations: rank-deficient and LoS channels. Actually, LoS channel may be viewed as a more serious rank-deficient channel.
5.1 Rank-deficient channel
In the section , as shown in Fig. 3, a rank-deficient channel is used. Here, BS transmits message to users via a cluster of scatters (CoS). The CoS has at least $K$ independent scattering points. The two channels from BS to CoS and IRS are assume to be full-rank. All channels from IRS and CoS to all users are assumed to be only LoS and they are rank-one channel, which is an extremely low-rank channel. System parameters are as follows: $M=512$, $N=60$, $Q=4$, $a=10^{-2.5}$. The positions of BS and IRS are (0m, 0m, 10m) and (80m, 20m, 20m), respectively. The six users are located at (100m, 0m, 0m), (85m, 5m, 0m), (95m, -5m, 0m), (85m, -10m, 0m), (90m, -20m, 0m) and (90m, -30m, 0m).
@@ 5.1 Rank-deficient channel
Fig. 4 displays sum-rate versus x-axis value of position of IRS $X$ in rank-deficient channel for the method of TLL-MMSE where $P_{I}=1$W, $P_{T}=0.7$W, $\sigma^{2}=-120$dBm and $\sigma_{z}^{2}=-40$dBm. As the x-axis value of IRS position varies from 50m to 150m by fixing its remaining axes. It can be observed from Fig. 4 that the sum-rate of the proposed TLL-MMSE with IRS has a substantial improvement over no IRS because of the doubled DoF. It is clear from this figure that the sum-rate first ascends monotonically, reaches its peak value, and descends monotonically. Thus, in this interval, the sum-rate can be viewed as a concave function of the x-axis value. Observing their maximum sum-rate values with IRS doubles those of no IRS due to the proved fact in Section 3: the number of DoFs with IRS in rank-deficient channel is twice that of no IRS.
@@
Fig. 5 shows the tendency curve of sum-rate versus noise power $\sigma^{2}$ at IRS of the proposed TLL-MMSE for different numbers of users, where $P_{I}=0.01$W, $P_{T}=5$W and $\sigma^{2}=\sigma_{z}^{2}$. As the noise power $\sigma^{2}$ raises from -140dBm to -60dBm, the sum-rate decreases gradually. This means that the noise level at active IRS will have a dominant influence on the sum-rate performance. To achieve a high-performance the system sum-rate, an active IRS with low noise is crucial.
@@
5.2 LoS channel
In the section all channels are assumed in LoS channel. Such a scenario is a typical rank-deficient channel with each channel being extremely low-rank, ie.rank-one. System parameters for this part are set as follows: $\sigma^{2}=\sigma_{z}^{2}=-40$dBm, $P_{T}=30$ dBm,$P_{I}=10$dBm, where $P_{T}$ presents total transmit power from BS. $M=16$, $Q=2$. The positions of BS and IRS are (0m, 0m, 10m) and (80m, 20m, 20m), respectively. The positions of four users are (100m, 0m, 0m), (100m, 10m, 0m), (90m, -20m, 0m) and (110m, -40m, 0m), respectively.
@@ 5.2 LoS channel
Fig. 8 illustrates the sum-rate versus the x-axis value of position of IRS $X$ when $N=16$. Its x-axis value changes from 50m to 150m by fixing the remaining axes of IRS. Similar to Fig. 4, the sum-rate function of $X$ is also a concave function over the x-axis range with the help of IRS. Obviously, there is at least one optimal position of IRS to achieve the maximum DoF.
[…]
@@ 6 Conclusions
In this paper, the achievable optimal DoF, and obtained sum-rate gains of an active IRS-aided MU-MIMO wireless communication network are investigated in LoS and rank-deficient channels, respectively. Specifically, we have derived the DoF upper bounds of IRS-aided single-user MIMO by using matrix inequalities and also extend the corresponding results to multi-user scenarios. In the rank-deficient channel, introducing IRS will make a dramatic DoF enhancement, i.e., doubling the value of DoF. To examine the rate gains achieved by enhanced DoF, subsequently, three beamforming methods that NSP-MTP-MRP, SO-MMSE and TLL-MMSE are designed to achieve the maximum DoF. Finally, from simulation results, in a rank-deficient channel, the devised TLL-MMSE aided by IRS may achieve a doubled sum-rate over no IRS, even in the small-scale or medium-scale scenarios, for example $N$ = 60. This is mainly due to the doubled increment in DoF induced by IRS. In LoS channel, the rate of the combined method NSP-MTP-MRP is more than twice that of no IRS for the SU-MIMO scenario in LoS channel when the size of IRS goes to large-scale. In summary, the immanent DoF enhancement ability of IRS in rank-deficient channel makes it become one of the most promising techniques for the future generation wireless communication networks like 6G, particularly a kind of low-rank-channel scenarios such as satellite communication, Space communications, UAV communications, and marine communications.

### v3

@@ 4.1 Proposed TLL-MMSE in rank-deficient channel
$+\sum_{j=1}^{K}{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}+\sigma^{2}_{z}\mathbf{I}_{k})\mathbf{U}_{k}$
@@
$+\sum_{j=1}^{K}{\color[rgb]{0,0,0}\sigma^{2}_{irs,k}}\mathbf{G}_{jk}^{H}\mathbf{\Theta}_{j}\mathbf{\Theta}_{j}^{H}\mathbf{G}_{jk}$
@@ 4.2 Proposed NSP-MTP-MRP in LoS channel
$+\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{n}+\mathbf{u}_{0}^{H}{\color[rgb]{0,0,0}\mathbf{z}_{1}},$ (65)
where $\mathbf{v}_{0}$ and $\mathbf{u}_{0}^{H}$ are transmit and receive beamforming vectors of user 111 from reflect channel respectively, and $\mathbf{\Theta}$ is the reflection coefficient matrix of IRS. And $\mathbf{v}_{k}$ and $\mathbf{u}_{k}^{H}$ are transmit and receive beamforming vectors of user $k$ from direct channel.
@@
$+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}\|\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\|^{2}+\sigma^{2}_{z}\|\mathbf{u}_{0}^{H}\|^{2}.$ (66)
@@
where $A_{0}=\sigma^{2}_{z}\|\mathbf{u}_{0}^{H}\|^{2}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}\|\mathbf{u}_{0}^{H}\mathbf{G}_{1}^{H}\mathbf{\Theta}\|^{2}$.
@@
$+\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{n}+\mathbf{u}_{k}^{H}{\color[rgb]{0,0,0}\mathbf{z}_{k}}.$ (69)
@@
where $A_{k}={\color[rgb]{0,0,0}\sigma^{2}_{irs}}\|\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\|^{2}+\sigma^{2}_{z}\|\mathbf{u}_{k}^{H}\|^{2}$.
@@
$P_{I}$ $=\hat{\mu}^{2}(\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}),$ (84)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{I}}{\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}}}.$ (85)
@@ 4.3 Proposed SO-MMSE in LoS channel
$P_{I}$ $=\hat{\mu}^{2}(\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}),$ (87)
@@
$\hat{\mu}$ $=\sqrt{\frac{P_{I}}{\|\text{diag}(\mathbf{H}_{BI}\mathbf{v}_{0})\hat{\boldsymbol{\theta}}^{H}\|^{2}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}}}.$ (88)
@@
$+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}\mathbf{u}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k}\mathbf{u}_{k}+\sigma^{2}_{z}\mathbf{u}_{k}^{H}\mathbf{u}_{k},$ (90)
@@
$\frac{\partial f(\mathbf{u}_{k})}{\partial\mathbf{u}_{k}}$ $=2({\color[rgb]{0,0,0}\sigma^{2}_{irs}}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k}+\sigma^{2}_{z}\mathbf{I}_{k}+\mathbf{H}_{dk}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{H}_{dk}^{H}$
@@
$\mathbf{H}_{dk}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{H}_{dk}^{H}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{k})^{-1}(\mathbf{H}_{dk}\mathbf{v}_{k}).$ (94)
@@
$\mathbf{H}_{d1}\mathbf{v}_{1}\mathbf{v}_{1}^{H}\mathbf{H}_{d1}^{H}+{\color[rgb]{0,0,0}\sigma^{2}_{irs}}\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G}_{1})^{-1}(\mathbf{G}_{1}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{v}_{0}).$ (95)
@@ 5 Simulation Results
In the following, simulation results are presented to verify the DoF gains achieved by IRS and assessed the corresponding sum-rate performance of the designed three methods. Below, simulations are conducted in two different situations: LoS plus Rayleigh fading channels and LoS plus LoS channels, where LoS plus Rayleigh fading channels represent that the channel BS-user is LoS channel and the BS-IRS-user channel is Rayleigh fading channel, and LoS plus LoS channels represent that the channel BS-user and the BS-IRS-user channel are LoS channels.
5.1 LoS plus Rayleigh fading channels
@@
In the section , Fig. 3 is the system model for the scenario of LoS plus Rayleigh fading channels. The system parameters are set as follows: $M=256$, $N=60$, $K=4$, $Q_{k}=4$, $P_{T}=40$ dBm, $P_{I}=30$ dBm, $\sigma_{u}^{2}=-120$ dBm, and $\sigma_{irs}^{2}=-90$ dBm. BS, IRS and $K$ users are respectively located at (100 m, 20 m, 0 m), (100 m, -20 m, 0 m), (100 m, 60 m, 0 m), and (100 m, -60 m, 0 m).
@@
Fig. 4 shows the curves of sum-rate versus IRS power $P_{I}$ for LoS plus Rayleigh fading channels under different numbers of users. Compared to no IRS, the sum-rate with IRS is higher especially in the high-power region. As $P_{I}$ increases from 0 dBm to 40 dBm, the sum-rate with IRS grows in general.
@@
Fig. 5 plots the curves of sum-rate versus power $P_{I}$ at IRS for the proposed TLL-MMSE method and existing NSP-ZF-PA method in [31]. From Fig. 5, it can be seen that the sum-rate of both methods increase as $P_{I}$ grows. This indicates that increasing the IRS power effectively improves the communication sum-rate performance. In addition, the proposed TLL-MMSE method performs better than NSP-ZF-PA as $P_{I}$ varies from 0 dBm to 40 dBm in terms of sum-rate. In particular, when $K=4$ and $P_{I}$ = 40 dBm, the sum-rate achieved by the proposed TLL-MMSE method is about 5 bit/s/Hz higher than that of NSP-ZF-PA method. It indicates that the the proposed TLL-MMSE method can effectively suppress the interference and improve the resource utilization.
[…]
Fig. 6 displays the curves of sum-rate versus transmit power $P_{T}$ at BS for LoS plus Rayleigh fading channels with different numbers of users when $\sigma_{u}^{2}=-50$ dBm and $\sigma_{irs}^{2}=-70$ dBm. As $P_{T}$ varies from 0 dBm to 40 dBm, the sum-rate increases as $P_{T}$ grows in general. Compared to no IRS, the sum-rate with IRS is much higher given a fixed $P_{T}$.
[…]
In accordance with (6), the SNRs corresponding to the different BS transmit power $P_{T}$ as shown in the axis-x of Fig. 6 can be calculated and plotted in Fig. 7. In this figure, as the transmit power at BS $P_{T}$ increases, the SNR per stream grows gradually. Compared to no IRS, the SNR with IRS is higher. In other words, IRS may improve the receive SNR at user. When $P_{T}=$ 25 dBm, the average SNRs at users are up to 30.9 dB (no IRS ) and 63 dB (with IRS), respectively. Clearly, the two SNR values are high. For a better explanation, the SNRs and sum-rates corresponding to the different BS transmit power when $K$=2 are listed in Table 1.
[…]
P_T (dBm) 0 5 10 15 20 25 30 35 40
no IRS SNR (dB) 10.9 15.9 20.9 25.9 30.9 35.9 40.9 45.9 50.9
no IRS sum-rate (bit/s/Hz) 7.4 10.6 13.9 17.2 20.5 23.8 27.1 30.5 33.8
with IRS SNR (dB) 38.0 43.0 48.0 53.0 58.0 63.0 68.0 73.0 78.0
with IRS sum-rate (bit/s/Hz) 24.7 28.0 31.3 34.6 38.0 41.3 44.6 48.0 51.2
Fig. 8 presents that the curves of sum-rate versus $\sigma_{irs}^{2}$ at IRS for the proposed TLL-MMSE method. From Fig. 8, it is seen that with the aid of IRS, the sum-rate of the multi-user MIMO network decreases gradually as $\sigma_{irs}^{2}$ grows from -120 dBm to -60 dBm. When $\sigma_{irs}^{2}$ is small, the IRS-assisted multi-user MIMO network can attain a higher sum-rate compared to the network without IRS. In the high $\sigma_{irs}^{2}$ region, the role of IRS gradually weakens, which make the sum-rate gap between the multi-user MIMO network with IRS and without IRS gradually decrease, which means that $\sigma_{irs}^{2}$ at active IRS has a dominant influence on the sum-rate performance. From Table 2, it is also easy to know that the sum-rate of the IRS-assisted multi-user MIMO network is about twice that of the network without IRS when $\sigma_{n}^{2}\leq-90$ dBm. When $K$ = 2 and $\sigma_{n}^{2}=-120$ dBm, the sum-rate of the IRS-assisted multi-user MIMO network even reaches up to 2.54 times that of the network without IRS.
[…]
$\sigma_{n}^{2}$ (dBm) K=4 K=2
$With~IRS$ $No~IRS$ ratio $With~IRS$ $No~IRS$ ratio
-120 49.5 19.5 2.54 49.5 19.5 2.54
-110 48.7 19.5 2.50 48.7 19.5 2.50
-100 46.4 19.5 2.38 46.4 19.5 2.38
-90 44.0 19.5 2.26 44.0 19.5 2.26
5.2 LoS plus LoS channels
[…]
@@ 5.2 LoS plus LoS channels
Fig. 12 illustrates the sum-rate versus the x-axis value of position of IRS $X$ when $N=16$. Its x-axis value changes from 50m to 150m by fixing the remaining axes of IRS. The sum-rate function of $X$ is a concave function over the x-axis range with the help of IRS. Obviously, there is at least one optimal position of IRS to achieve the maximum DoF. It is also easy to know that the conclusion applies to both low-SNR and high-SNR conditions.
@@ 6 Conclusions
In this paper, the achievable optimal DoF, and obtained sum-rate gains of an active IRS-aided MU-MIMO wireless communication network are investigated in LoS and low-rank channels, respectively. Specifically, we have derived the DoF upper bounds of IRS-aided single-user MIMO by using matrix inequalities and also extend the corresponding results to multi-user scenarios. In the low-rank channel, introducing IRS will make a dramatic DoF enhancement, i.e., doubling the upper bound of DoF. To examine the rate gains achieved by enhanced DoF, subsequently, three beamforming methods that NSP-MTP-MRP, SO-MMSE and TLL-MMSE are designed to achieve the maximum DoF. Finally, from simulation results, in LoS plus Rayleigh fading channels, the devised TLL-MMSE aided by IRS may achieve a 2.54 times sum-rate over no IRS, even in the small-scale or medium-scale scenarios, for example $N$ = 60. This is mainly due to the doubled increment in DoF induced by IRS. In LoS plus LoS channels, the rate of the combined method NSP-MTP-MRP is more than twice that of no IRS for the SU-MIMO scenario in LoS plus LoS channels when the size of IRS goes to large-scale. In summary, the immanent DoF enhancement ability of IRS in low-rank channel makes it become one of the most promising techniques for the future generation wireless communication networks like 6G, particularly a kind of low-rank-channel scenarios such as satellite communication, space communications, UAV communications, and marine communications.
@@ 5 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grant U22A2002, and by the Hainan Province Science and Technology Special Fund under Grant ZDYF2024GXJS292; in part by the Scientific Research Fund Project of Hainan University under Grant KYQD(ZR)-21008; in part by the Collaborative Innovation Center of Information Technology, Hainan University, under Grant XTCX2022XXC07; in part by the National Key Research and Development Program of China under Grant 2018YFB180110.
@@
  • [31] K. Yang, R. Dong, W. Gao, F. Shu, W. Shi, Y. Wang, Multi-Stream Transmission for Directional Modulation Network via Distributed Muli-UAV-Aided Multi-Active-RS, lEEE Open Joumal of the Communications Society, early access.
  • [32] D. Xu, X. Yu, D. W. Kwan Ng, and R. Schober, Resource allocation for active IRS-assisted multiuser communication systems, in Proc. 55th Asilomar Conf. Signals Syst. Comput., Pacific Grove, CA, USA, 2021, pp. 113–119.
  • [33] M. Munawar and K. Lee, Low-Complexity Adaptive Selection Beamforming for IRS-Assisted Single-User Wireless Networks, IEEE Trans. Veh. Technol., vol. 72, no. 4, pp. 5458-5462, 2023.